# CLUE: A Clinical Language Understanding Evaluation for LLMs

### v3

@@ Abstract
Large Language Models (LLMs) are expected to significantly contribute to patient care, diagnostics, and administrative processes. Emerging biomedical LLMs aim to address healthcare-specific challenges, including privacy demands and computational constraints. Assessing the models’ suitability for this sensitive application area is of the utmost importance. However, evaluation has primarily been limited to non-clinical tasks, which do not reflect the complexity of practical clinical applications. To fill this gap, we present the Clinical Language Understanding Evaluation (CLUE), a benchmark tailored to evaluate LLMs on clinical tasks. CLUE includes six tasks to test the practical applicability of LLMs in complex healthcare settings. Our evaluation includes a total of $2525$ LLMs. In contrast to previous evaluations, CLUE shows a decrease in performance for nine out of twelve biomedical models. Our benchmark represents a step towards a standardized approach to evaluating and developing LLMs in healthcare to align future model development with the real-world needs of clinical application. We open-source all evaluation scripts and datasets for future research: https://github.com/TIO-IKIM/CLUE
CLUE: A Clinical Language Understanding Evaluation for LLMs
@@ 1 Introduction
Figure 1: The proposed CLUE Benchmark consists of six clinical tasks ranging from answering patient inquiries to multi-document knowledge extraction and simplification.
@@
The introduction of Large Language Models (LLMs) into the healthcare sector marks the beginning of a transformative period for medical practitioners, promising significant advancements in the quality and efficiency of patient care Moor et al. (2023). In clinical research settings, these advances are driven by open-source LLMs, specifically the development of biomedical LLMs (Luo et al., 2023; Chen et al., 2023; Labrak et al., 2024). These models are tailored to the medical sector to deliver improved performance with fewer parameters than their general-domain counterparts.
Despite the promise these models hold, their evaluation has predominantly focused on constructed medical questions, such as exam quizzes (Han et al., 2023; Toma et al., 2023; Chen et al., 2023; Labrak et al., 2024; Gururajan et al., 2024; Griot et al., 2024). While indicative of the models’ medical knowledge, these assessments do not adequately reflect the complexity and diversity of clinical settings they are intended for.
Moreover, the number of released general and biomedical LLMs is immense and biomedical evaluations settings are often not overlapping or even contradict each other Touvron et al. (2023a); Han et al. (2023); Jiang et al. (2023); Touvron et al. (2023b); Toma et al. (2023); Chen et al. (2023); Li et al. (2023); Tunstall et al. (2023); Jiang et al. (2024); Labrak et al. (2024); Gururajan et al. (2024); Griot et al. (2024); Christophe et al. (2024); Abdin et al. (2024). This rapid evolution poses challenges in determining the most appropriate LLMs for clinical applications at any given moment, including assessing whether the latest general LLMs surpass the performance of previous domain-specific LLMs. To clarify this highly dynamic landscape, it is essential to implement ongoing evaluations using a standardized benchmark.
@@
Catering to this demand, our study introduces the Clinical Language Understanding Evaluation (CLUE), a novel benchmark specifically designed to assess LLMs in the context of practical downstream tasks in healthcare in a comparable setting. CLUE is based on clinical documents and provides a comprehensive framework for evaluating the performance of LLMs. The benchmark consists of six clinical tasks that probe a diverse set of model abilities, including medical reasoning, information extraction, simplification, coding, and summarization, offering a robust challenge to test the essential capabilities of these models in a clinical setting. This allows us to analyze the strengths and shortcomings of $2525$ models. We include open-source models to address typical demands for data privacy regulations in clinical environments Liu et al. (2023); Wang et al. (2023).
In summary, our initiative to develop CLUE and conduct a thorough evaluation of both biomedical and general-domain LLMs aims to bridge the current knowledge gap in testing these technologies in healthcare. By challenging models with a wide set of clinical tasks, our benchmark seeks to pave the way for more informed decisions in the deployment of LLMs in medical settings, ultimately enhancing patient care and streamlining the workload of healthcare professionals.
@@ 2.1 Large Language Models in the clinical domain
With the transformer architecture (Vaswani et al., 2017), automatic language processing gained noticeable traction (Hossain et al., 2019; Wolf et al., 2019), which has led to the development of initial foundation models for not only general language processing (Devlin et al., 2018; Conneau and Lample, 2019; Raffel et al., 2020) but also the specialized medical context (Liu et al., 2021; Yan et al., 2022; Rasmy et al., 2021). This progression has only been accelerated since the reveal of ChatGPT in late 2022 as the potential use for dynamically usable LLMs has been spread through the public. Commercial models such as Palm Chowdhery et al. (2022) and GPT4 (Achiam et al., 2023) seem to deliver great performances for various tasks that at first glance might rival specialists (Yunxiang et al., 2023; Nori et al., 2023; Lee et al., 2023; Thirunavukarasu et al., 2023), but on more thorough investigation show noticeable discrepancies (Lee et al., 2023). Shortly after the release of commercial models, a surge of open-source alternatives like Llama (Touvron et al., 2023a; Wu et al., 2023), Alpaca (Taori et al., 2023) and Mistral (Jiang et al., 2023) followed them. To deliver better performance on medical tasks, these models were pre-trained on medical data (Han et al., 2023; Peng et al., 2023a, b; Labrak et al., 2024). The specific evaluation settings, such as whether a model is fine-tuned on downstream tasks, how many examples are given, and techniques like chain-of-thought prompting, often vary, making the results impossible to compare. CLUE addresses this by evaluating models in a uniform setting on clinically relevant tasks.
@@ 2.2 Medical evaluation of LLMs
Evaluation of medical LLMs has predominantly focused on datasets comprising multiple choice exam questions, such as MMLU (Hendrycks et al., 2021), MEDQA (Jin et al., 2021), MedMCQA (Pal et al., 2022), and synthetic questions generated from PubMed articles (Jin et al., 2019). These tasks are a good way to test the representation of medical knowledge in models and, to some extent, their medical reasoning capabilities. However, they fail to address several critical aspects necessary for effective real-world medical applications, such as the comprehension of complex patient cases from semi-structured documents with medical jargon and abbreviations, the reformulation of information for tasks like summarization, the translation between lay and specialist language, and the answering of open-ended questions. Furthermore, recent work has reported severe issues with data contamination of LLM evaluation (Magar and Schwartz, 2022; Shi et al., 2024), specifically with regards to MMLU (Deng et al., 2023), and proposed solutions to mitigate these issues (Li et al., 2024; Deng et al., 2023).
This evaluation paradigm highlights a notable gap in that none of the models were assessed using clinical text datasets. Despite the requirement for extensive medical knowledge and reasoning in addressing these quizzes, the nature of the texts vastly differs from that of clinical documents. Clinical documents are characterized by their irregular structure, frequent jargon, typing errors, and abbreviations that vary across medical specialties and regions, and their typically greater length than quiz questions, leading to a less homogeneous domain of texts.
DR.BENCH (Gao et al., 2023) is a benchmark focusing on clinical documents but predominantly consists of traditional encoder model tasks, such as extractive question-answering, sequence tagging, and text classification, rendering it less suitable for a comprehensive evaluation of LLMs. Kweon et al. (2024) introduced a multiple-choice question-answering dataset derived from MIMIC-IV Electronic Health Records (EHR), and Adams et al. (2024) released another MCQA dataset based on fictive reports concerning constructed patient cases. However, both studies refrained from evaluating biomedical LLMs on their datasets. CLUE combines appropriate existing tasks with novel tasks for a standardized clinical evaluation of general-domain and biomedical models.
3 CLUE Benchmark
CLUE includes six clinical document tasks, which encompass information extraction, summarization, clinical reasoning, simplification, and coding. These tasks are designed to better assess the practical applications of LLMs in clinical settings. Table 1 in Appendix summarizes the characteristics of these tasks. We provide prompt examples for each task in Figures 9, 10, 11, 12, 13, and 14 in Appendix B.3.
@@ 3.1.1 Dataset Generation
In the first phase, we split the text into sentences (for details, see A.1). In the second phase, we prompted the LLM to generate question-answer pairs from the list of sentences for each document. We instruct the LLM to ground each answer in one or more sentences so we can check the quality manually. After completing the first two phases, we have a list of question-answer pairs. In the third phase, these candidates are manually reviewed, and high-quality examples are selected based on the following criteria:
@@
Figure 8 in Appendix A.1 lists some examples of problematic pairs and selected pairs.
As a last step, we remove the discharge letters from the original documents to pair the questions and answers with the resulting documents.
@@ 4 Experimental setup
Figure 4: The average level 1 and level 2 scores. Models that share the same base are displayed in the same color.
We evaluated $2525$ language models, including general domain and specialized biomedical models. Since some biomedical models were not instruction-tuned, we also included general domain base models for a fair comparison. Our evaluation aims to (1) rank current openly available models, (2) assess whether biomedical models or general domain instruction-tuned models are more suitable for specific clinical scenarios, and (3) measure the effects of continuous biomedical training.
@@ 4.1 General-domain models
We include both the instruction-tuned and base checkpoints of Llama2-7B, Llama2-70B Touvron et al. (2023b), Mistral-7B-v0.1 Jiang et al. (2023), Meta-Llama-3-8B-Instruct and Meta-Llama-3-70B-Instruct AI@Meta (2024) since they serve as a basis for several biomedically fine-tuned models. Additionally, we include zephyr-7b-beta Tunstall et al. (2023), Mistral-7B-Instruct-v0.2 Jiang et al. (2023), Phi-3-mini-128k-instruct Abdin et al. (2024), Mixtral-8x7B and Mixtral-8x22B Jiang et al. (2024) due to their general popularity.
Since the context length of Llama2-based models is insufficient to read all MIMIC discharge summaries and LongHealth documents, we split the evaluation into two levels: The first one with input sizes of less than 1.000 words on average and level two with more than 1.000 on average with long documents. Table 1 in Appendix A shows the average number of words per input for all tasks.
4.2 Biomedical Models
Table 2 in B.2 shows the biomedical LLMs we evaluate in our experiments. Meditron-7B and 70B (Chen et al., 2023) are continuously pretrained from Llama-2-7B Touvron et al. (2023a). The pretraining dataset consists of 48.1B tokens (Chen et al., 2023). In contrast to Meditron, internistai/base-7b-v0.2 (Griot et al., 2024) was only trained on 2.3B tokens with a focus on data quality based on Mistral-7B-v0.1 (Jiang et al., 2023). BioMistral (Labrak et al., 2024) is another continuously pre-trained model that was based on Mistral-7B-Instruct-v0.1. They use weight-merging algorithms such as DARE (Yu et al., 2023) to combine the weights of their biomedical model and the original instruction-tuned model. Toma et al. (2023) released ClinicalCamel, a language model based on Llama-2-70B. A more recent model was published by Gururajan et al. (2024) and is based on Meta’s Llama3 (AI@Meta, 2024). They employed Direct Preference Optimization (DPO) and supervised fine-tuning (SFT) to train instruction-tuned models. A similar fine-tuning procedure was done for Llama3-OpenBioLLM-8B and Llama3-OpenBioLLM-70B (Ankit Pal, 2024). However, no information on the training process has been provided yet. Finally, the models JSL-Med-Sft-Llama-3-8B John Snow Labs (2024a), JSL-MedLlama-3-8B-v1.0 John Snow Labs (2024b) and JSL-MedLlama-3-8B-v2.0 John Snow Labs (2024c) also fine-tuned Llama3 variants with very limited available information.
Figure 5: The relative difference between biomedical models and their base model.
@@ 5 Results
We present the benchmark results in Figures 4 and 5. Detailed scores for all models are in Table 3 of Appendix C.
Figure 4 shows that overall the general-domain models Meta-Llama-3-70B-Instruct, Mixtral-8x22B-Instruct-0.1 and Meta-Llama-3-8B-Instruct score the highest. The biomedical model Llama3-OpenBioLLM-70B is also among these models, however clearly behind its base model Meta-Llama-3-70B-Instruct. The lower left quadrant is mainly filled with biomedical models, with some scoring below the baseline.
For well-performing models, the difference between level 1 scores is generally smaller than the difference between level 2 scores. This indicates the expressiveness of LongHealth and the newly introduced tasks MeDiSumQA and MeDiSumCode.
5.1 Effect of biomedical fine-tuning
For many biomedical models, there is a clear decline in performance, which is visualized in Figure 5. While models based on Mistral-7B-v0.1 improve up to $5.525.52$ points, models based on Meta-Llama-3 show a very clear decrease with drops of up to $31.3731.37$ points. The remaining models, Meditron-7B, Meditron-70B, and ClinicalCamel-70B, are also not able to improve the original Llama2-70B performance, while the instruction-tuned general domain Llama2-70B very clearly outperforms the biomedical models. (see Appendix C).
@@ 5.2 Error Analysis
Major contributions to the performance drop of biomedical models are task 3 of LongHealth and the MeDiSumCode valid code scores. Since LongHealth Task 3 measures how often a model returns no answer if the requested information is not in the input document, it indicates the frequency of hallucinations. Similarly, the valid code scores in MeDiSumCode show models often invent ICD-10 codes. Models that have a low valid code score often start to count up the numbers in ICD codes instead of actually predicting valid codes. Examples of this behavior are listed in Appendix C.1. LongHealth task 3 reveals a drastic change from Meta-Llama-3-8B-Instruct, which scored $56.2556.25$, to Llama3-OpenBioLLM-8B, which only scored $1.551.55$. Similarly, Llama3-OpenBioLLM-70B also performs poorly compared to Meta-Llama-3-70B-Instruct on this task.

### v4

# Does Biomedical Training Lead to Better Medical Performance?

@@ Abstract
Large Language Models (LLMs) are expected to significantly contribute to patient care, diagnostics, and administrative processes. Emerging biomedical LLMs aim to address healthcare-specific challenges, including privacy demands and computational constraints. Assessing the models’ suitability for this sensitive application area is of the utmost importance. However, biomedical training has not been systematically evaluated on medical tasks. This study investigates the effect of biomedical training in the context of six practical medical tasks evaluating $2525$ models. In contrast to previous evaluations, our results reveal a performance decline in nine out of twelve biomedical models after fine-tuning, particularly on tasks involving hallucinations, ICD10 coding, and instruction adherence. General-domain models like Meta-Llama-3.1-70B-Instruct outperformed their biomedical counterparts, indicating a trade-off between domain-specific fine-tuning and general medical task performance. We open-source all evaluation scripts and datasets to support further research in this critical area.
Does Biomedical Training Lead to Better Medical Performance?
@@ 1 Introduction
Figure 1: Our evaluation consists of six medical tasks with increasing complexity ranging from answering patient inquiries to multi-document knowledge extraction and simplification.
@@
The introduction of Large Language Models (LLMs) into the healthcare sector marks the beginning of a transformative period for medical practitioners, promising significant advancements in the quality and efficiency of patient care Moor et al. (2023). In medicine, a major factor for these advances are open-source LLMs, specifically biomedical LLMs (Luo et al., 2023; Chen et al., 2023; Labrak et al., 2024). These models are tailored to the medical sector to deliver improved performance with fewer parameters than their general-domain counterparts.
Despite the promise these models hold, their evaluation has predominantly focused on constructed medical questions, such as exam quizzes (Han et al., 2023; Toma et al., 2023; Chen et al., 2023; Labrak et al., 2024; Gururajan et al., 2024; Griot et al., 2024). While multiple-choice quizzes offer a controlled environment for assessment, real-world medical tasks are often open-ended, involve incomplete information that the model must recognize to avoid generating inaccurate responses, and require processing lengthy, unstructured inputs that challenge the model’s ability to organize and extract relevant data efficiently. This raises the question of whether biomedical training methods are also improving LLM performance on more practical medical tasks.
Dataset Samples Mean Words Metrics Focus
MedNLI 1425 20.81 Accuracy Clinical reasoning
MeQSum 1000 60.77 R-L, R-1, R-2, BERT F1 Summarization
LongHealth 400 5536.82 Accuracy Information extraction
Problem Summary 237 123.5 R-L, R-1, R-2, BERT F1, UMLS F1 Information extraction
Table 1: An overview of the characteristics of the tasks.
Moreover, the number of released general and biomedical LLMs is immense and biomedical evaluations settings are often not overlapping or even contradict each other Touvron et al. (2023a); Han et al. (2023); Jiang et al. (2023); Touvron et al. (2023b); Toma et al. (2023); Chen et al. (2023); Li et al. (2023); Tunstall et al. (2023); Jiang et al. (2024); Labrak et al. (2024); Gururajan et al. (2024); Griot et al. (2024); Christophe et al. (2024b); Abdin et al. (2024). Different settings are often used for evaluation, which makes it even more difficult to assess the general effects of biomedical training of models.
@@
Our study addresses these challenges by evaluating LLMs before and after biomedical training on a set of practical downstream tasks in healthcare in a comparable setting. We evaluate on six tasks that probe a diverse set of model abilities, including medical reasoning, information extraction, simplification, ICD coding, and summarization, offering a robust challenge to test a range of essential capabilities of these models in a medical setting. Figure 1 summarizes the key features of these tasks. This allows us to analyze the strengths and shortcomings of $1212$ models before and after biomedical training.
Despite biomedically trained models often outperforming their general domain counterparts on multiple-choice quizzes, our study concludes that these results do not always translate to practical tasks. In fact, multiple biomedical LLMs perform worse on practical tasks than general models. This can be attributed to increased hallucinations and reduced accuracy and stability, suggesting that while specialized training enhances domain-specific knowledge, it may also introduce overfitting or bias, leading to unreliable outputs in complex, real-world medical scenarios. By challenging models with a wide set of medical tasks, our evaluation seeks to pave the way for more informed decisions in the deployment of LLMs in medical settings, ultimately enhancing patient care and streamlining the workload of healthcare professionals.
@@ 2.1 Large Language Models in the clinical domain
The development of biomedical Large Language Models (LLMs) has gained considerable momentum in recent years, driven by the need for more specialized tools in the healthcare sector. This rapid growth is evident in the emergence of both commercial models, such as Med-PaLM (Singhal et al., 2023) and MedGemini (Saab et al., 2024), and open-source alternatives like Meditron (Chen et al., 2023), Biomistral (Labrak et al., 2024), Internist.ai (Griot et al., 2024), and Med42 (Christophe et al., 2024b).
Biomedical LLMs have shown a marked improvement over their general-domain counterparts, particularly in multiple-choice question-answering (MCQA) tasks, where they benefit from their domain-specific training. In some cases, these models have even challenged the performance of leading commercial tools like GPT-4 (Christophe et al., 2024a), highlighting their potential in handling specialized medical knowledge.
However, while biomedical fine-tuning has led to improved performance on MCQA tasks, it remains unclear whether these gains translate to practical medical tasks.
@@ 2.2 Medical evaluation of LLMs
Evaluation of medical LLMs has primarily centered on datasets composed of multiple-choice questions, such as MMLU (Hendrycks et al., 2021), MEDQA (Jin et al., 2021), MedMCQA (Pal et al., 2022), and synthetic questions from PubMed articles (Jin et al., 2019). While these tasks assess medical knowledge representation and some reasoning abilities, they overlook crucial aspects needed for real-world applications, such as understanding complex patient cases, translating between lay and specialist language, and addressing open-ended questions. This evaluation approach reveals a gap—models aren’t tested on clinical text datasets. Unlike exam questions, clinical documents are irregularly structured, rife with jargon, errors, abbreviations, and generally longer, making them a more heterogeneous domain.
Several studies have evaluated LLMs in medical contexts. Gao et al. (2023) assessed realistic medical tasks, though they did not include few-shot evaluations or modern LLM architectures. Kweon et al. (2024) and Adams et al. (2024) focused on realistic input documents like EHRs, without specifically evaluating biomedical LLMs. Chen et al. (2024) concentrated on biomedical models with MCQA tasks but did not explore practical tasks or compare performance before and after biomedical training. Liu et al. (2024) covered a broad range of tasks, including practical and biomedical models, though a systematic comparison of pre- and post-training performance was not conducted, and advanced models like Med42-v2 (Christophe et al., 2024a) and Meditron3 (OpenMeditron, 2024) were not considered.
Notably, none of these studies reported negative effects of biomedical training. In contrast, our study systematically compares $1212$ biomedical LLMs with general-domain counterparts on practical medical tasks and finds that biomedical training can lead to increased hallucinations and decreased accuracy, highlighting potential drawbacks of specialized training.
3 Evaluation Tasks
Our evaluation includes six clinical document tasks, encompassing information extraction, summarization, clinical reasoning, simplification, and coding. Two of these are introduced in this study to better assess the practical applications of LLMs in clinical settings. Table 1 summarizes the characteristics of these tasks. We provide prompt examples for each task in Figures 8, 9, 10, 11, 12, and 13 in Appendix B.3.
@@ 3.1.1 Dataset Generation
In the first phase, we split the text into sentences (for details, see A.1). In the second phase, we prompted the LLM to generate question-answer pairs from the list of sentences for each document. We instruct the LLM to ground each answer in one or more sentences so we can check the quality manually. After completing the first two phases, we have a list of question-answer pairs. In the third phase, these candidates are manually reviewed by a medical expert, and high-quality examples are selected based on the following criteria:
@@
Figure 7 in Appendix A.1 lists some examples of problematic pairs and selected pairs. As a last step, we remove the discharge letters from the original documents to pair the questions and answers with the resulting documents.
@@ 4 Experimental setup
Figure 4: The average level 1 and level 2 scores. Models that share the same base are displayed in the same color. Arrows indicate the performance difference to the base model after biomedical training.
We evaluated $2525$ language models, including biomedically trained models, their base models, and additional general-domain models as reference. Our evaluation aims to (1) measure the effects of continuous biomedical training, (2) assess whether biomedical models or general domain models are more suitable for specific medical scenarios, and (3) rank current openly available models.
4.1 Biomedical Models
Table 4 in B.2 shows the biomedical LLMs we evaluate in our experiments. Meditron-7B and 70B (Chen et al., 2023) are continuously pretrained from Llama-2-7B Touvron et al. (2023a). The pretraining dataset consists of 48.1B tokens (Chen et al., 2023). In contrast to Meditron, internistai/base-7b-v0.2 (Griot et al., 2024) was only trained on 2.3B tokens with a focus on data quality based on Mistral-7B-v0.1 (Jiang et al., 2023). BioMistral (Labrak et al., 2024) is another continuously pre-trained model that was based on Mistral-7B-Instruct-v0.1. They use weight-merging algorithms such as DARE (Yu et al., 2023) to combine the weights of their biomedical model and the original instruction-tuned model. A more recent model was published by Gururajan et al. (2024) and is based on Meta’s Llama3 (AI@Meta, 2024). They employed Direct Preference Optimization (DPO) (Rafailov et al., 2024) and supervised fine-tuning (SFT) to train instruction-tuned models. A similar fine-tuning procedure was used for Llama3-OpenBioLLM-8B and Llama3-OpenBioLLM-70B (Ankit Pal, 2024). However, no information on the training process has been provided yet. Med42-Llama3-8B and Med42-Llama3-70B (Christophe et al., 2024b) were trained on a mixture of various medical datasets, as well as general domain datasets using and iterative alignment approach (Christophe et al., 2024b). Finally, Meditron3-8B and Meditron3-70B (OpenMeditron, 2024) were trained on a similar dataset as the first Meditron. However, this generation was trained from instruction-tuned Llama3.1 models (Dubey et al., 2024) and the data also included instructions.
@@ 4.2 General-domain models
Apart from the previously mentioned models we also evaluated zephyr-7b-beta Tunstall et al. (2023), Mistral-7B-Instruct-v0.2 Jiang et al. (2023), Phi-3-mini-128k-instruct Abdin et al. (2024), Mixtral-8x7B and Mixtral-8x22B Jiang et al. (2024) to compare the performance of biomedical models with general-domain models.
4.3 Levels
Since the context length of Llama2-based models is insufficient to read all MIMIC discharge summaries and LongHealth documents, we split the evaluation into two levels: The first one with input sizes of less than 1.000 words on average and level two with more than 1.000 on average with long documents. Table 1 shows the average number of words per input for all tasks. We evaluate models with a context size of less than 8k tokens only on level 1 tasks.
@@ 5 Results
We present the benchmark results in Figure 4, with detailed scores for all models available in Table 5 in Appendix C.
Figure 4 illustrates that models often perform lower after biomedical training, especially when the respective base model performs well. Notably, this applies to all models based on Llama3 and Llama3.1 except Llama3-Med42-8B which performs better on simple level 1 tasks, but worse on level 2 tasks.
Models based on Mistral-7B, which is the lowest-scoring general-domain model, show a performance improvement. However, the improvement of BioMistral-7B is only marginal, and internistai/base-7b-v0.2 only improves on level 2 tasks. Additionally, they do not demonstrate a clear advantage over zephyr-7b-beta, another general-domain model based on Mistral-7B.
Generally, Llama3 and Llama3.1-based biomedical models show a clearer performance drop on level 2 tasks than level 1 tasks, indicating the expressiveness of LongHealth and the newly introduced tasks, MeDiSumQA and MeDiSumCode.
LH Task3 Valid Codes
BioMistral-7B +4.15 +17.26
BioMistral-7B-DARE +0.95 +18.79
internistai/base-7b-v0.2 +45.55 +16.32
Llama3-OpenBioLLM-8B -40.05 -10.77
Llama3-Med42-8B -12.7 -6.8
Llama3-Aloe-8B-Alpha -22.55 -17.09
Llama3-OpenBioLLM-70B -28.80 -20.29
Llama3-Med42-70B -79.15 -15.39
Meditron3-8B -52.15 -49.19
Meditron3-70B -54.6 -4.76
Table 2: Mistral-7B-v0.1 , Meta-Llama-3-(8B/70B) and Meta-Llama-3.1-(8B/70B) based models on LongHealth task 3 and percentage of valid ICD-10 codes in MeDiSumCode
@@ 5.1 Error Analysis
Major contributions to the performance drop of biomedical models are task 3 of LongHealth and the MeDiSumCode valid code scores. Table 2 shows the performance difference between biomedical and their base models on these metrics. All models based on Mistral-7B show a clear performance improvement, while all Llama3-based models perform worse.
Since LongHealth task 3 measures how often a model returns no answer if the requested information is not in the input document, it indicates the frequency of hallucinations. Similarly, the valid code scores in MeDiSumCode show models often invent ICD-10 codes. Models that have a low valid code score often start to count up the numbers in ICD codes instead of predicting valid codes. Examples of this behavior are listed in Appendix C.1. LongHealth task 3 reveals a drastic change from Meta-Llama-3-8B-Instruct, which scored $56.2556.25$, to Llama3-OpenBioLLM-8B, which only scored $1.551.55$. Similarly, Llama3-OpenBioLLM-70B also performs poorly compared to Meta-Llama-3-70B-Instruct on this task.